### v1

# MHE under parametric uncertainty - Robust state estimation without informative data

Abstract
In this paper, we study state estimation for general nonlinear systems with unknown parameters and persistent process and measurement noise. In particular, we are interested in stability properties of the state estimate in the absence of persistency of excitation (PE). With a simple academic example, we show that existing moving horizon estimation (MHE) approaches as well as classical adaptive observers can result in diverging state estimates in the absence of PE, even if the noise is small. We propose a novel MHE formulation involving a regularization based on a constant prior estimate of the unknown system parameters. Only assuming the existence of a stable estimator, we prove that the proposed MHE results in practically robustly stable state estimates even in the absence of PE. We discuss the relation of the proposed MHE formulation to state-of-the-art results from MHE, adaptive estimation, and functional estimation. The properties of the proposed MHE approach are illustrated with a numerical example of a car with unknown tire friction parameters.
{IEEEkeywords}
Estimation; Moving horizon estimation; Parametric uncertainty; Absence of persistency of excitation
1 Introduction
State estimation is of high importance for control, system monitoring, and fault detection in many applications such as power systems [1], battery management systems [2], manufacturing systems [3], or biomedical systems [4]. Obtaining an accurate estimate of the state in general requires access to a model which accurately describes the underlying system dynamics. Identifying the parameters of such a model exactly is, however, difficult, e.g., due to the nonlinear nature of the models. In addition, the parameters are often slowly varying, e.g., due to aging of components, and therefore practical applications require some form of online adaptation of the model parameters. This renders the simultaneous estimation of states and parameters of nonlinear systems from noisy output measurements a problem of high practical relevance, compare, e.g, [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, the resulting estimation problem is nonlinear and non-convex even in the simple case of linear systems, see, e.g., [16, 17], and thus challenging to solve. Furthermore, unique identification of model parameters requires sufficiently informative data, typically through a so-called persistency of excitation (PE) condition (cf. [18, 19, 20, 21]), even if noise-free state measurements are available. Reliable parameter estimation is not possible in the absence of PE or, more generally, if the excitation is small compared to the noise, i.e., the signal to noise ratio is small. For general nonlinear systems or over-parametrized models, ensuring PE is typically impossible. For linear systems, PE can be ensured using additional excitation [20], however, this requires active deterioration of the regulation performance. In this paper, we present an estimation method with guaranteed robust stability for general nonlinear systems with unknown parameters subject to process and measurement noise, without relying on a PE assumption.
Related work
Adaptive observers are classical methods for online state estimation under parametric uncertainty [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Assuming PE, convergence of state and parameter estimates resulting from such an adaptive observer can be shown for linear systems [22, 23, 24, 25, 26] and selected classes of nonlinear systems such as SISO systems in observable canonical form [27] or linearly parameterized SISO systems [28]. Additionally, robust stability under process and measurements noise was established for adaptive observers for linear systems in [30, 29]. Under a less restrictive initial excitation (IE) condition, e.g., dynamic regressor extension and mixing [33, 34] or concurrent learning [34] lead to stable estimates in linear systems without noise. Assuming bounded noise and IE, robustness of the estimation under process and measurement noise was analyzed for linear systems in [35] by employing a switching mechanism in the parameter estimator. In the absence of PE, boundedness of state and parameter estimation errors in the noise-free case, i.e., without process and measurement noise, were shown for linear systems [32, 34] and a special class of nonlinear systems [30, 31], where the nonlinearity in the systems dynamics depends only on inputs and measurements. Ensuring robustness of state estimates under parametric uncertainty in the absence of PE is in general an open problem, which is challenging to address since PE is virtually necessary for robustness of parameter estimates [36]. Overall, existing adaptive observer designs are mostly applicable to linear systems or restrictive classes of nonlinear systems, and typically lack robustness guarantees in the absence of PE.
More recently, joint state and parameter estimation for nonlinear systems has been addressed using optimization-based approaches, see, e.g., [37, 38] for the case of linear systems. Moving horizon estimation (MHE) is a promising optimization-based technique for estimation in general nonlinear systems subject to process and measurement noise [39, 40, 41]. In particular, uniform [^1] Uniform in the control input. robust stability of MHE can be established with a sufficiently large horizon under a general detectability condition in the form of incremental input/output-to-state-stability ($\delta$-IOSS) [42, 40]. Compared to classical estimation approaches, MHE requires us to solve a finite-horizon optimization problem online, which lead to the development of efficient solution strategies [10, 43, 44, 45]. Joint state and parameter estimation in MHE is commonly addressed by introducing an augmented state containing both the original system states and the unknown parameters [46, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Such MHE approaches are employed in different applications reaching from autonomous driving and friction estimation [5, 6], estimation in vibrating structures [7, 8], chemical processes [9, 10, 11, 12], biological processes [13, 14], to induction motors [15]. Uniform detectability ($\delta$-IOSS) of the augmented state consisting of the system state and parameters in general requires a uniform PE condition  [47, 48]. Consequently, the (robust) stability properties of these MHE schemes deteriorate arbitrarily if the excitation becomes small during online operation. Additionally, applying an MHE for joint state and parameter estimation in the absence of PE can result in numerical issues. While this has been addressed by modifications of the prior weighting in the MHE objective, compare [48, 49], the resulting MHE schemes do not provide any stability guarantees in the absence of PE. Using recent results from functional estimation [50], stability of the state estimate can be ensured even if the augmented state is not detectable. However, this requires the use of a full information estimator (FIE), i.e., an optimization-based estimator with unbounded horizon. The design of finite horizon MHE approaches with robust stability properties in the absence of PE for general nonlinear systems with parametric uncertainty and process and measurement noise remains an open question.
Contribution
In this paper, we derive an MHE formulation for general nonlinear systems with parametric uncertainty subject to process and measurement noise which ensures reliable state estimation in the absence of PE, and consequently without unique identifiability of the parameters. With a simple academic example, we show that standard MHE schemes for joint state and parameter estimation can lead to parameter drift in the absence of PE and thus lack fundamental robust stability properties (Section 3). We also demonstrate that classical adaptive observers have the same limitation (Section 5.3). To overcome this, we propose a novel MHE design by replacing the standard prior weighting of the parameter estimate with a simple regularization (Section 4). We show practical robust stability of the resulting state estimation error under the proposed MHE approach. Robust stability implies that the estimation error is proportional to the process and measurement noise. The stability is practical with respect to (w.r.t.) a constant involving the bias of the parameter regularization, which can be made arbitrarily small by increasing the horizon. To establish robust stability we only require a weak detectability condition on the system: There exists an estimator satisfying the desired robust stability property. Notably, satisfaction of this assumption does not require detectability of the augmented system state and hence does not impose a restrictive PE condition. The presented theoretical analysis relies on recent techniques developed in the context of MHE [40] and generalizations thereof to functional estimation [50]. We discuss the relation of the proposed MHE approach to classical MHE schemes and existing prior weightings to handle lack of PE, connect the proposed ”regularization” to classical tools from adaptive control and estimation, and discuss the relations to functional estimation (Section 5). Finally, we present a numerical example of a car with unknown tire friction parameters and in the absence of PE (Section 6).
Notation
Let the non-negative real numbers be denoted by $\mathbb{R}_{\geq 0}$, the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a$ for some $a\in\mathbb{R}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for some $a,b\in\mathbb{R}$ with $a\leq b$ by $\mathbb{I}_{[a,b]}$. Let $\|x\|$ denote the Euclidean norm of the vector $x\in\mathbb{R}^{n}$. A positive definite (semi-definite) matrix $P=P^{\top}$ is denoted as $P\succ 0$ ($P\succeq 0$). The quadratic norm with respect to a positive definite matrix $Q=Q^{\top}$ is denoted by $\|x\|_{Q}^{2}=x^{\top}Qx$, and the minimal and maximal eigenvalues of $Q$ are denoted by $\lambda_{\min}(Q)$ and $\lambda_{\max}(Q)$, respectively. The maximum generalized eigenvalue of positive definite matrices $A=A^{\top}$ and $B=B^{\top}$ is denoted as $\lambda_{\max}(A,B)$, i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The $i$-th element of a vector $x\in\mathbb{R}^{n}$ is denoted as $[x]_{i}$ for $i\in\mathbb{I}_{[1,n]}$.
2 Setup and Preliminaries
In this section, we formulate the considered state estimation problem, introduce the desired stability property, and state the necessary assumption to ensure the existence of an appropriate estimator.
2.1 Problem Formulation
We consider a nonlinear discrete-time system
$x_{t+1}$ $=f(x_{t},u_{t},w_{t},\theta),$ (1a)
$y_{t}$ $=h(x_{t},u_{t},w_{t},\theta),$ (1b)
where $t\in\mathbb{I}_{\geq 0}$ is the discrete time step, $x_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$ the system state, $u_{t}\in\mathbb{U}=\mathbb{R}^{m}$ the control input, $w_{t}\in\mathbb{R}^{q}$ the process and measurement noise, $\theta\in\mathbb{R}^{n_{\theta}}$ a constant but unknown parameter, and $y_{t}\in\mathbb{Y}=\mathbb{R}^{p}$ the measurement output. Note that $w_{t}$ appears in the dynamics (1a) and measurement model (1b) and hence can also model separate process disturbances and measurement noise.
We consider the general case where we have additional system information in the form of constraints
$x_{t}\in\mathbb{X},\ w_{t}\in\mathbb{W},\ \forall t\in\mathbb{I}_{\geq 0},\ \theta\in\Theta,$ (2)
with $\mathbb{X}\subseteq\mathbb{R}^{n_{x}}$, $\mathbb{W}\subseteq\mathbb{R}^{q}$, and $\Theta\subseteq\mathbb{R}^{n_{\theta}}$. Note that this is not restrictive, since we can always choose $\mathbb{X}=\mathbb{R}^{n_{x}}$, $\mathbb{W}=\mathbb{R}^{q}$, and $\Theta=\mathbb{R}^{n_{\theta}}$ in case no additional system knowledge is available. Constraints of the form (2) can be used to include additional physical prior knowledge, e.g., non-negativity of certain states, or bounds on noise and parameters, into a state estimator, compare [39, Sec. 4.3.3]. In the following, we assume that the true trajectory of system (2.1) always satisfies the constraints (2).
Given estimates $\bar{x}_{0}$ and $\bar{\theta}_{0}$ of the initial state $x_{0}$ and the unknown parameter $\theta$, respectively, the objective is to obtain an estimate $\hat{x}_{t}$ of the current system state $x_{t}$ at each time step $t\in\mathbb{I}_{\geq 0}$. The considered problem is formalized by the following definition of a state estimator.
Definition 1 (State estimator, adapted from [51, Def. 2.2])
A state estimator is a sequence of functions $\Psi_{t}:\mathbb{X}\times\Theta\times\mathbb{W}^{t}\times\mathbb{Y}^{t}\times\rightarrow\mathbb{X}$ for $t\in\mathbb{I}_{\geq 0}$ to compute estimates of (1a) at each time step $t$ as
$\hat{x}_{t}=\Psi_{t}\left(\bar{x}_{0},\bar{\theta}_{0},\{\bar{w}_{j}\}_{j=0}^{t-1},\{\bar{y}_{j}\}_{j=0}^{t-1}\right),$ (3)
where $\bar{w}_{j}$ and $\bar{y}_{j}$ are estimates of the noise $w_{j}$ and measurements $y_{j}$ at time step $j$, respectively.
In a practical application, the noise estimates $\bar{w}_{t}$ are typically chosen to be zero and the output measurement $\bar{y}_{j}$ equal to the noisy output measurement $y_{j}$ according to (1b). The aim is to design a state estimator according to Definition 1 for system (2.1) which satisfies the following stability notion.
Definition 2 (Incremental input-to-output practical stability)
A state estimator according to Definition 1 is exponentially incrementally input-to-output practically stable ($\delta$-IOpS) with respect to some constant $\epsilon\geq 0$ if there exist $\lambda_{1},\lambda_{2},\lambda_{3}\in[0,1)$ and $C_{1},C_{2},C_{3}>0$ such that
$\|x_{t}-\hat{x}_{t}\|\leq$ $\max\left\{C_{1}\lambda_{1}^{t}\left\|\begin{bmatrix}x_{0}\\ \theta\end{bmatrix}-\begin{bmatrix}\bar{x}_{0}\\ \bar{\theta}_{0}\end{bmatrix}\right\|\right.,$ (8)
$\max_{j\in\mathbb{I}_{[0,t-1]}}C_{2}\lambda_{2}^{t-j-1}\|w_{j}-\bar{w}_{j}\|,$ (9)
[…]
A state estimator according to Definition 1 is exponentially incrementally input-to-output stable ($\delta$-IOS) if additionally $\epsilon=0$.
This stability notion is an extension of the one defined in [50, Def. 3] in the context of functional estimation, compare also [52, Def. 6] for classical input-to-state practical stability. In contrast to classical stability, Inequality (9) only provides a bound on the estimation error of the state $x_{t}$, but not the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$. Note that in case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, and (9) is satisfied with $\epsilon=0$, Definition 2 corresponds to a standard robust stability notion considered in MHE, compare [40, Def. 2]. To design an $\delta$-IOpS state estimator, we require the following assumption toward the dynamics of the considered system.
Assumption 1 (Existence of stable state estimator)
There exists an exponentially $\delta$-IOS state estimator according to Definition 2.
Note that only the existence of such an estimator is assumed, but it does not need to be known or have finite complexity. In Section 5.5 below, we establish the existence of such an estimator for a common class of systems.
2.2 Detectability
In the following, we introduce an incremental input/output-to-output-stability ($\delta$-IOOS) Lyapunov function as introduced in [50], which characterizes existence of a stable estimator (Assumption 1).
Definition 3 ($\delta$-IOOS Lyapunov function [50, Def. 6])
A function $W_{\delta}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{\theta}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}_{\geq 0}$ is an exponential $\delta$-IOOS Lyapunov function if there exist $P_{1},\ P_{2,x},\ P_{2,\theta}\succ 0$ and $Q,\ R\succeq 0$ such that
$\|x-\tilde{x}\|_{P_{1}}^{2}\leq W_{\delta}(x,\tilde{x},\theta,\tilde{\theta})\leq\|x-\tilde{x}\|_{P_{2,x}}^{2}+\|\theta-\tilde{\theta}\|_{P_{2,\theta}}^{2},$ (13a)
$W_{\delta}\left(x^{+},\tilde{x}^{+},\theta,\tilde{\theta}\right)\leq\eta W_{\delta}(x,\tilde{x},\theta,\tilde{\theta})$
$\hskip 113.81102pt+\|w-\tilde{w}\|_{Q}^{2}+\|y-\tilde{y}\|_{R}^{2},$ (13b)
for all $\{x,u,w,y\}\ \in\ \mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{Y}$, $\{\tilde{x},u,\tilde{w},\tilde{y}\}\ \in\ \mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{Y}$, $\theta\in\Theta$, and $\tilde{\theta}\in\Theta$, where $x^{+}=f(x,u,w,\theta)$, $\tilde{x}^{+}=f(\tilde{x},u,\tilde{w},\tilde{\theta})$, $y=h(x,u,w,\theta)$, and $\tilde{y}=h(\tilde{x},u,\tilde{w},\tilde{\theta})$.
In the following proposition, we establish that Assumption 1 is a necessary and sufficient condition for the existence of a $\delta$-IOOS Lyapunov function according to Definition 3.
Proposition 1
Assumption 1 holds if and only if system (2.1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3.
Proof 2.1.
The considered problem is a special case of functional estimation studied in [50] with system state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ and corresponding functional output $z_{t}=\phi(x_{t},\theta)=x_{t}$ to be estimated. Resulting from [50, Cor. 8], system (2.1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3 if and only if the system is exponentially $\delta$-IOOS according to [50, Def. 4]. Additionally, it was shown in [50] that exponential $\delta$-IOOS [^2] Both [50, Prop. 5] and [50, Cor. 10] consider a more general asymptotic form of $\delta$-IOOS, but the proof naturally extends to the case of exponential $\delta$-IOOS. is both necessary (cf. [50, Prop. 5]) and sufficient (cf. [50, Cor. 10]) for Assumption 1 to be satisfied. Consequently, Assumption 1 is necessary and sufficient for the existence of an exponential $\delta$-IOOS Lyapunov function.
Existing stability results for MHE instead use a $\delta$-IOSS Lyapunov function, which would require $W_{\delta}$ to be lower bounded in the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$, cf. [51][40]. In [47], it was shown that this detectability condition also holds for the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ if $x_{t}$ is detectable and an additional a PE condition holds. In contrast, the detectability condition considered here does not depend on a PE condition, and thus holds uniformly for any sequence of inputs.
Note that we restrict our analysis to an exponential-like version of $\delta$-IOOS here. In [50, Thm. 9] a more general asymptotic version of $\delta$-IOOS was used to establish stability of a FIE approach for functional estimation. Notably, also without parametric uncertainty, establishing stability for an MHE scheme with finite horizon requires a more restrictive (exponential) condition, compare [41, Sec. 5.5.3] and [40, Rem. 4].
2.3 Standard MHE for State and Parameter Estimation
In the following, we introduce a standard MHE formulation to jointly estimate states and parameters, which is widely applied, see, e.g., [46, 5, 6, 7, 8, 9, 10, 12, 11, 13, 15, 14]. The MHE approach at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past noises and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ within a window of length $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state and parameter estimates, $\hat{x}_{t-M_{t}}$ and $\hat{\theta}_{t-M_{t}}$, respectively. Thereby, the MHE scheme optimizes over a sequence of $M_{t}$ disturbance estimates $\hat{w}_{\cdot|t}=\{\hat{w}_{j|t}\}_{j=t-M_{t}}^{t-1}$, and the initial state and the parameter estimates, $\hat{x}_{t-M_{t}|t}$ and $\hat{\theta}_{t}$, respectively. Using the system model (2.1), the disturbance sequence, initial state estimate, and parameter estimate define a sequence of state $\{\hat{x}_{j|t}\}_{j=t-M_{t}}^{t}$ and output $\{\hat{y}_{j|t}\}_{j=t-M_{t}}^{t-1}$ estimates. The cost function is chosen as
$V_{\mathrm{MHE}}^{\mathrm{x},\theta}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)=$
$\quad 2\eta^{M_{t}}\left(\|\hat{x}_{t-M_{t}|t}-\hat{x}_{t-M_{t}}\|_{P_{\mathrm{x}}}^{2}+\|\hat{\theta}_{t}-\hat{\theta}_{t-M_{t}}\|_{P_{\theta}}^{2}\right)$ (14)
$\quad+2\sum_{j=1}^{M_{t}}\eta^{j-1}\left(\|\hat{w}_{t-j|t}-\bar{w}_{t-j}\|_{Q}^{2}+\|\hat{y}_{t-j|t}-\bar{y}_{t-j}\|_{R}^{2}\right),$
with a possible discount factor $\eta\in[0,1]$ and weighting matrices $P_{\mathrm{x}},\ P_{\theta},\ R,\ Q\succ 0$. The two first terms in (14) involving the past state and parameter estimates, $\hat{x}_{t-M_{t}}$ and $\hat{\theta}_{t-M_{t}}$, are called prior weighting on the state and parameter, respectively. The MHE objective (14) is most commonly chosen without discounting, i.e., with $\eta=1$ [46, 5, 6, 7, 8, 9, 10, 12, 11, 13, 15, 14]. The state and parameter estimates at time step $t$ are then obtained by solving the following nonlinear program (NLP)
$\min_{\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{\theta}_{t}}$ $V_{\mathrm{MHE}}^{\mathrm{x},\theta}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)$ (15a)
s.t. $\hat{x}_{j+1|t}=f(\hat{x}_{j|t},u_{j},\hat{w}_{j|t},\hat{\theta}_{t}),\ j\in\mathbb{I}_{[t-M_{t},t-1]},$ (15b)
$\hat{y}_{j|t}=h(\hat{x}_{j|t},u_{j},\hat{w}_{j|t},\hat{\theta}_{t}),\ j\in\mathbb{I}_{[t-M_{t},t-1]},$ (15c)
$\hat{w}_{j|t}\in\mathbb{W},\ \hat{y}_{j|t}\in\mathbb{Y},\ j\in\mathbb{I}_{[t-M_{t},t-1]},$ (15d)
$\hat{x}_{j|t}\in\mathbb{X},\ j\in\mathbb{I}_{[t-M_{t},t]},$ (15e)
$\hat{\theta}_{t}\in\Theta.$ (15f)
We denote a (not necessarily unique) minimizer of (2.3) as $\hat{x}_{t-M_{t}|t}^{*}$, $*\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $*\hat{x}_{\cdot|t}^{*}$ and $*\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state and parameter estimates at time step $t$ are denoted as $\hat{x}_{t}=\hat{x}_{t|t}^{*}$ and $\hat{\theta}_{t}=\hat{\theta}_{t}^{*}$, respectively. The MHE scheme (2.3) is applied in receding horizon fashion. Thereby, the current state and parameter estimates are obtained at each time step $t$ by solving the (finite-horizon) MHE problem (2.3) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weightings based on the state and parameter estimates at time step $t-M_{t}$.
Stability of such an MHE scheme can be established using tools from [40] if the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is detectable, the horizon length $M$ is chosen sufficiently large, and if $\eta\in[0,1)$. In [47] it was shown that the underlying detectability conditions for the augmented state holds under a PE condition.
3 Motivating Example
Figure 1: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach for state and parameter estimation (2.3).
In this section, we provide an academic example to demonstrate that standard MHE formulations (Section 2.3) lack fundamental robust stability properties in the absence of PE. The presented example is inspired by [53, Sec. V], which demonstrates parameter drift in adaptive filtering. Consider the system
$x_{t+1}=$ $ax_{t}+u_{t}+[w_{t}]_{1},$ (16a)
$y_{t}=$ $\begin{bmatrix}x_{t}+[w_{t}]_{2}\\ \theta x_{t}+[w_{t}]_{3}\end{bmatrix},$ (16d)
with $a=0.99$, $[w_{t}]_{1}=0$, $[w_{t}]_{2}=-0.1$, and $[w_{t}]_{3}=0.1$ for all $t\in\mathbb{I}_{\geq 0}$. Thereby, the constant values of $[w_{t}]_{2}$ and $[w_{t}]_{3}$ can be interpreted as sensor biases, which often occur in practice. Further consider $x_{0}=1$, $\bar{x}_{0}=1$, $\theta=1$, and $\bar{\theta}_{0}=1.1$, and the MHE formulation (2.3) with $M=20$, $\eta=0.99$, $Q=I_{3}$, $R=I_{2}$, $P_{\mathrm{x}}=10$, $P_{\theta}=10$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. Note that
$x_{t}=a\left([y_{t-1}]_{1}-[w_{t-1}]_{2}\right)+u_{t-1}+[w_{t-1}]_{1},$
implies that a stable state estimator exists and hence that system (3) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3 with any $\eta\in(0,1)$ (compare Proposition 1). The input $u_{t}$ is chosen as
$u_{t}=\begin{cases}1&t\mod 10000=0,\\ 0&\text{else}.\end{cases}$
The input is zero over long periods of time and the state asymptotically converges to zero. Hence, the ratio of the excitation signal compared to the noise becomes very small. As shown in Figure 1, the parameter diverges for such a standard MHE formulation. Consequently, whenever the output contains informative data (when the input is equal to $1$), the drifted parameter estimate leads to an increasing state estimation error. This phenomena is known as bursting in classical adaptive control [54, 55] and is depicted in the zoom-in in Figure 1. Thus, robust convergence of the resulting state and parameter estimate is not guaranteed. Note that this problem equally occurs if an MHE scheme (2.3) without discounting ($\eta=1$) is applied. In the remainder of the paper, we propose a novel MHE formulation which overcomes this issue by using a different prior weighting, which allows us to establish stability in the state estimates even in the absence of PE.
4 Moving Horizon Estimation in the Absence of Persistency of Excitation
In this section, we introduce an MHE formulation (Section 4.1) to estimate the state of a nonlinear dynamical system subject to parametric uncertainty in the absence of PE. Furthermore, we prove that the proposed MHE is $\delta$-IOpS according to Definition 2 (Section 4.2), which ensures that the stability issue in the absence of PE, as discussed in the previous section, is overcome (Section 4.3).
4.1 Moving Horizon Estimation Formulation
Equivalently to the standard MHE formulation stated in Section 2.3, our proposed MHE at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past disturbances and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ in a window of length $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state estimate $\hat{x}_{t-M_{t}}$. However, instead of using the past parameter estimate $\hat{\theta}_{t-M_{t}}$, the proposed MHE formulation relies on the prior $\bar{\theta}_{0}$ of the unknown parameter. The objective of the optimization-based state estimation problem is chosen as
$V_{\mathrm{MHE}}^{\mathrm{x}}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)=$
$\quad 2\eta^{M_{t}}\left(\|\hat{x}_{t-M_{t}|t}-\hat{x}_{t-M_{t}}\|_{P_{2,\mathrm{x}}}^{2}+\|\hat{\theta}_{t}-\bar{\theta}_{0}\|_{P_{2,\theta}}^{2}\right)$ (17)
$\quad+2\sum_{j=1}^{M_{t}}\eta^{j-1}\left(\|\hat{w}_{t-j|t}-\bar{w}_{t-j}\|_{Q}^{2}+\|\hat{y}_{t-j|t}-\bar{y}_{t-j}\|_{R}^{2}\right),$
where $\eta\in[0,1)$. The discount factor $\eta$ and weighting matrices $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, and $Q$ are chosen based on the exponential $\delta$-IOOS Lyapuonv function according to Definition 3, which is guaranteed to exist due to Assumption 1 and Proposition 1. In Remark 4.6 below, we discuss how to design the cost (17) in a practical setting where a $\delta$-IOOS Lyapunov function is not known. The state estimate at time step $t$ is then obtained by solving the following NLP
$\min_{\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{\theta}_{t}}$ $V_{\mathrm{MHE}}^{x}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)$ (18a)
[…]
We denote a (non-unique) minimizer [^3] We assume that a minimizer to (2.3) always exists. Existence of such a minimizer can be ensured under mild assumptions, e.g., if $P_{2,\mathrm{x}},P_{2,\theta},Q,R\succ 0$, or the sets $\mathbb{W},\mathbb{Y},\mathbb{X},\Theta$ are compact, cf. [39, App. A.11]. of (4.1) as $\hat{x}_{t-M_{t}|t}^{*}$, $*\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $*\hat{x}_{\cdot|t}^{*}$ and $*\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state estimate at time step $t$ is denoted as
$\hat{x}_{t}=\hat{x}_{t|t}^{*},$ (19)
and the corresponding estimation error as
$\hat{e}_{t}=\hat{x}_{t}-x_{t}.$ (20)
The MHE estimator (4.1) is then applied in a receding horizon fashion. Thereby, the state estimate (19) at each time step $t$ is obtained by solving the optimization problem (4.1) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weighting based on the state estimate at time step $t-M_{t}$.
4.2 Theoretical Analysis
In the following, we establish that the MHE (4.1) is an exponentially $\delta$-IOpS estimator according to Definition 2 provided the horizon $M$ is chosen larger than some lower bound.
Theorem 4.2 (MHE is $\delta$-IOpS).
Let Assumption 1 hold and suppose the horizon $M$ satisfies
$\rho^{M}\coloneqq 4\eta^{M}\lambda_{\max}(P_{2,\mathrm{x}},P_{1})<1,$ (21)
with $\rho\in[0,1)$. Then, the MHE estimator (4.1) is exponentially $\delta$-IOpS according to Definition 2, i.e, (9) holds for all $t\in\mathbb{I}_{\geq 0}$ with (uniform) constants $C_{1},\ C_{2},\ C_{3}>0$, and
ϵ=italic-$\epsilon=$ $\ \frac{4}{\sqrt{1-\rho}}\sqrt{\frac{\lambda_{\max}(P_{2,\theta})}{\lambda_{\min}(P_{1})}}\sqrt{\rho}^{M}\|\bar{\theta}_{0}-\theta\|,$ (22)
$\lambda_{1}=$ $\ \sqrt{\rho},\ \lambda_{2}=\lambda_{3}=\sqrt[4]{\rho}.$
The proof of Theorem 4.2 can be found in Appendix .1. Note that the bound (9) with (22) resulting from Theorem 4.2 implies that for $M$ going to infinity, the MHE (4.1) is an exponentially $\delta$-IOS estimator according to Definition 2. It follows that FIE, i.e., Problem (4.1) with $M_{t}=t$, is exponentially $\delta$-IOS.
Corollary 4.3 (FIE is $\delta$-IOS).
Let Assumption 1 hold. Then, the FIE estimator (4.1) with $M_{t}=t$ is exponentially $\delta$-IOS according to Definition 2.
Proof 4.4.
The proof follows directly from (44) by setting $M=t$.
The above corollary shows that the results from [50, Thm. 9] are recovered, compare also our discussion regarding functional estimation in Section 5.4 below.
Remark 4.5 (Known parameter).
In case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, it follows directly from Theorem 4.2, and in particular (22), that the estimation error (20) resulting from the proposed MHE (4.1) is robustly globally exponentially stable (with $\epsilon=0$) according to [56, Def. 1].
Remark 4.6 (Parameterization of MHE objective).
Choosing $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, and $Q$ in (17) based on a $\delta$-IOOS Lyapunov function $W_{\delta}$ (Definition 3) might seem to limit the applicability of the proposed MHE design and theory since such a Lyapunov function is generally not known explicitly in practical applications. However, due to Assumption 1 and Proposition 1 some $W_{\delta}$ with $\tilde{P}_{2,\mathrm{x}}$, $\tilde{P}_{2,\theta}$, $\tilde{R}$, $\tilde{Q}\succ 0$ and $\tilde{\eta}\in[0,1)$ is guaranteed to exist. Due to scale invariance of $W_{\delta}$, the weighting matrices $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, $Q\succ 0$ in the objective (17) can be chosen arbitrarily, provided $\eta>\tilde{\eta}$, i.e., $01-\eta>$ 0 is sufficiently small, and the horizon $M$ is long enough (Condition (21) in Theorem 4.2). Compare also a similar discussion in [40, Rem. 1].
4.3 Motivating Example Revisited
Figure 2: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach with prior on $\hat{\theta}_{t-M_{t}}$ (2.3) and our proposed MHE approach with prior $\bar{\theta}_{0}$ (4.1).
In the following, we revisit the academic example system (3) discussed in Section 3 and apply the proposed MHE (4.1) to it. We choose the same values for $\bar{\theta}_{0}$, $M$, $\eta$, $Q$, $R$ as for the standard MHE in Section 3, with $P_{2,\mathrm{x}}=P_{\mathrm{x}}$, and $P_{2,\theta}=P_{\theta}$. The corresponding absolute state estimation and parameter errors are shown in Figure 2. For comparison, the estimation results from Section 3 using a standard MHE for state and parameter estimation (2.3) are shown again. As expected from the theoretical analysis (Theorem 4.2), the state estimation error resulting from the proposed MHE is significantly smaller. In particular, the parameter error remains uniformly bounded and the parameter converges to the prior $\bar{\theta}_{0}$ whenever the data is not informative enough.
5 Discussion
In this section, we discuss the relation of the MHE approach presented in Section 4 to various existing methods for state and parameter estimation from the literature.
5.1 Standard Moving Horizon Estimation
Standard MHE approaches for state and parameter estimation similar to the one described in Section 2.3 are widely applied [46, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Assuming PE, robust convergence of the joint state and parameter estimation error system under a standard MHE scheme can be established [47, Thm. 2]. However, as seen in Section 3, lack of PE can lead to parameter drift and unstable state estimates. We overcome this issue by replacing the standard prior weighting on the parameter with a regularization with respect to a (constant) prior parameter estimate $\bar{\theta}_{0}$. This ensures practical stability of the state estimates irrespective of PE. In case of sufficiently large PE, the standard MHE formulation ensures robust stability, while the proposed MHE may result in less accurate state estimates due to the bias with respect to the prior parameter. Hence, while the proposed modification improves stability properties independent of PE, it may also deteriorate the estimation performance if PE is sufficiently large. Deriving an MHE formulation that combines both properties remains an interesting question for future research.
5.2 Adaptations of Prior Weighting in Absence of PE
The prior weighting in the MHE objective (14) is often chosen such that the solution of the MHE problem is as close as possible to the infinite-horizon (FIE) solution. In the linear unconstrained case with quadratic cost and no unknown parameters, the prior weighting can be computed using a standard Kalman filter such that the MHE solution matches exactly to the one of FIE [57]. Thus, the prior weighting in many nonlinear MHE formulations is computed using an extended Kalman filter [58, 59, 60, 61, 10]. In case of joint state and parameter estimation in the absence of PE, such a prior weighting becomes singular/ill-conditioned. Thus a minimizer to the optimization problem may not exist, which leads to a lack of robustness both from a robust and statistical point of view [62]. In [49], a regularization of the prior weighting in connection with pseudo measurements was introduced, thus reducing the sensitivity to measurement outliers, compare also [48] for a similar approach. These mechanisms are comparable to forgetting factors in classical adaptive estimation and control [63]. However, in contrast to the proposed MHE scheme, the approaches in [48, 49] with adapted prior weightings lack (robust) stability guarantees in the absence of PE.
5.3 Adaptive Estimation
Figure 3: Academic example (Section 3): Absolute state and parameter estimation errors resulting from the adaptive observer proposed in [32] for both the noise-free case ($w=0$) and the case with measurement noise.
Classically, adaptive observers are applied for state estimation in case of parametric uncertainty. While there exist adaptive observers with guarantees in the absence of PE [30, 31, 32], those approaches in general lack robustness against process and measurement noise. In Figure 3, we show the application of the adaptive estimation approach [^4] Application of the adaptive observer from [32] requires to transform the system (3) to observable canonical form, consequently having a state of dimension $n=2$. We choose the parameters of the observer from [32] as $F=a\cdot I_{2}$ (with $a=0.99$ from the dynamics introduced in Section 3), $R=I_{2}$, and apply covariance resetting [32, Sec. 3.1] with $k_{0}=0.1$ and $k_{\min}=0.01$. from [32] to the academic example presented in Section 3, which is characterized by measurement noise and absence of PE. It is visible that the parameter estimate converges in the noise-free case (with $w=0$), resulting in small state estimation errors. In the noisy case, we observe a similar effect as for the MHE with prior on $\hat{\theta}_{t-M_{t}}$ as shown in Figure 2. In particular, the slowly diverging parameter estimates lead to large state estimation errors whenever there is enough excitation. In contrast, the proposed MHE approach is robustly stable (Definition 2) for any general nonlinear system whenever there exists any robustly stable estimator (Assumption 1). Hence, whenever it is in principle possible to design an adaptive observer with robust stability properties, then, Assumption 1 holds and the proposed MHE formulation yields robust stability properties. However, to the best of the author’s knowledge, existing adaptive observer designs do not have such properties for general nonlinear systems.
The problem of parameter drift and the related bursting phenomena is well known in the adaptive control and estimation literature [64, 55, 53, 54] and the interplay between PE and robustness remains actively researched [36, 33]. We propose a simple constructive solution to this problem by introducing a regularization cost on the parameter estimate w.r.t. a constant initial parameter estimate $\bar{\theta}_{0}$. As shown in [65], such a regularization can be related to the classical $\sigma$-modification, compare [66, Chap. 8][67], which ensures boundedness/robustness irrespective of PE.
5.4 Functional Estimation
Functional estimation [68, 69, 50] is applied to estimate a function of the state $z_{t}=\phi(x_{t})$. In [50, Thm. 9], it was shown that robustly stable estimates of $z_{t}$ can be obtained using an FIE if the system admits a stable estimator (Assumption 1). The state estimation problem considered here is a special case of a functional estimation problem for the augmented system with state $\tilde{x}_{t}=\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ with functional output $\tilde{z}_{t}=x_{t}$, compare [50, Rem. 1]. Since this functional estimation problem is nonlinear (even for linear systems), classical functional estimation approaches [68, 69] are not applicable. Corollary 4.3 recovers the FIE results for functional estimation from [50, Thm. 9]. In addition, such an FIE can even ensure (asymptotic) robust stability under a weaker asymptotic detectability condition [50, Thm. 9]. However, in contrast to the proposed MHE formulation with fixed horizon length, solving FIE is in general not computationally tractable.
The MHE design and theoretical analysis presented here show that it is possible to establish (practical) robust stability for functional estimation using MHE, but under certain (potentially restrictive) conditions. In particular, to establish (practical) robust stability of the estimates of the system state $x_{t}$, it is crucial that the (non-detectable) parameter $\theta$ is constant. Finding general conditions on the functional estimation problem which enable (practical) robust stability of MHE for functional estimation is an interesting topic of future research.
5.5 Verification of Assumption 1 in Practical Applications
Our theoretical results in Section 4.2 only require the existence of a robustly stable state estimator (Assumption 1). While this assumption is clearly necessary, it is nontrivial to verify whether it holds in practical applications. In the following, we show that the existence of a stable estimator (Assumption 1) can be established for a very common class of (mechanical) systems. In particular, consider a discrete-time system of the form
$\begin{bmatrix}p_{t+1}\\ v_{t+1}\\ y_{t}\end{bmatrix}=\begin{bmatrix}p_{t}+\Delta tA(p_{t})v_{t}\\ v_{t}+\Delta tf(p_{t},v_{t},u_{t},\theta)\\ p_{t}\end{bmatrix}+w_{t},$ (29)
with state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{2\cdot n_{\mathrm{p}}}$, control input $u_{t}\in\mathbb{R}^{m}$, parameter $\theta\in\mathbb{R}^{n_{\theta}}$, measurement $y_{t}\in\mathbb{R}^{n_{\mathrm{p}}}$, process and measurement noise $w_{t}\in\mathbb{R}^{3\cdot n_{\mathrm{p}}}$, sampling time $\Delta t$, and $A^{-1}(p)$ Lipschitz continuous. Such a model (29) can describe the behavior of many mechanical systems where the position $p$ can be measured, e.g., quadrotors [40, Sec. V.B], and cars [70], compare also the presented numerical example in Section 6 below.
In the noise-free case (i.e., with $w=0$), the state of system (29) can be exactly recovered given two past outputs (32), even without knowledge of the parameters. Hence, a robustly stable state estimator can be constructed, i.e., Assumption 1 is satisfied [^5] In the noise-free case, the states can be recovered as $p_{t}=y_{t}$, $v_{t}=\frac{1}{\Delta t}A^{-1}(y_{t})\left(y_{t+1}-y_{t}\right)$. Due to Lipschitz continuity of $A^{-1}$, the state estimation error is bounded proportional to the magnitude of the process and measurement noise, i.e., Equation (9) holds with arbitrarily small positive constants $\lambda_{1},\lambda_{2},\lambda_{3}>0$ and $\epsilon=0$. However, note that this estimator is not causal, which slightly deviates from Definition 1.. On the other hand, the augmented state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ of system (29) is in general not uniformly detectable, compare, e.g., the numerical example in Section 6 below.
6 Practical Application Example
In this section, we present a numerical example of state estimation under parametric uncertainty for a car in a realistic scenario. We show that a standard MHE approach for joint state and parameter estimation results in detrimental state estimates due to parameter drift, while the proposed MHE approach ensures robust state estimation. In addition, we show that the proposed adaptation allows for superior state estimation compared to an MHE with fixed parameter value and consequently a persistent parameter error.
6.1 State Estimation for Car
We consider a dynamic bicycle model with simplified Pacejka tire force model [71, 70]. The continuous-time system is governed by following differential equations
$\dot{x}_{\mathrm{p}}$ $=v_{\mathrm{x}}\cos(\psi)-v_{\mathrm{y}}\sin(\psi),$ (30a)
$\dot{y}_{\mathrm{p}}$ $=v_{\mathrm{x}}\sin(\psi)+v_{\mathrm{y}}\cos(\psi),$ (30b)
$\dot{\psi}$ $=\omega,$ (30c)
$\dot{v}_{\mathrm{x}}$ $=\frac{1}{m}\left(F_{\mathrm{x}}-F_{\mathrm{f}}\sin(\delta)+mv_{\mathrm{y}}\omega\right),$ (30d)
$\dot{v}_{\mathrm{y}}$ $=\frac{1}{m}\left(F_{\mathrm{r}}+F_{\mathrm{f}}\cos(\delta)-mv_{\mathrm{x}}\omega\right),$ (30e)
$\dot{\omega}$ $=\frac{1}{I_{\mathrm{z}}}\left(F_{\mathrm{f}}l_{\mathrm{f}}\cos(\delta)-F_{\mathrm{r}}l_{\mathrm{r}}\right),$ (30f)
Table 1: Parameters of miniature car from [70]
$l_{\mathrm{r}}$ $l_{\mathrm{f}}$ $m$ $I_{\mathrm{z}}$
$0.038$ $0.052$ $0.181$ $5.05\cdot 10^{-4}$
$B_{\mathrm{r}}$ $C_{\mathrm{r}}$ $D_{\mathrm{r}}$ $B_{\mathrm{f}}$ $C_{\mathrm{f}}$ $D_{\mathrm{f}}$
$8.5$ $1.45$ $1.0$ $5.2$ $1.5$ $0.65$
where $x_{\mathrm{p}}$, $y_{\mathrm{p}}$, $\psi$ are the x and y coordinates of the car position and the yaw angle in world frame, $\omega$ is the yaw rate, and $v_{\mathrm{x}}$ and $v_{\mathrm{y}}$ are the longitudinal and lateral velocities in body frame of the car. Further, $m$ is the mass of the car, $I_{\mathrm{z}}$ is the inertia along the $z$-axis, and $l_{\mathrm{f}}$ and $l_{\mathrm{r}}$ are the distance of the front and rear axis from the center of mass, respectively. The lateral tire forces $F_{\mathrm{f}}$ and $F_{\mathrm{r}}$ are modeled with the simplified Pacejka tire force model
$\alpha_{\mathrm{f}}$ $=\arctan\left(\frac{v_{\mathrm{y}}+\omega l_{\mathrm{f}}}{v_{\mathrm{x}}}\right)-\delta,$ (31a)
$\alpha_{\mathrm{r}}$ $=\arctan\left(\frac{v_{\mathrm{y}}-\omega l_{\mathrm{r}}}{v_{\mathrm{x}}}\right),$ (31b)
$F_{\mathrm{f}}$ $=D_{\mathrm{f}}\sin(C_{\mathrm{f}}\arctan(B_{\mathrm{f}}\alpha_{\mathrm{f}})),$ (31c)
$F_{\mathrm{r}}$ $=D_{\mathrm{r}}\sin(C_{\mathrm{r}}\arctan(B_{\mathrm{r}}\alpha_{\mathrm{r}})),$ (31d)
where $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$ are the front and rear slip angles, and $B_{\mathrm{f}},\ B_{\mathrm{r}},\ C_{\mathrm{f}},\ C_{\mathrm{r}},\ D_{\mathrm{f}}$, and $D_{\mathrm{r}}$ are the Pacejka tire model parameters. The control input is $u=\begin{bmatrix}\delta&F_{\mathrm{x}}\end{bmatrix}^{\top}\in\mathbb{R}^{2}$ and the overall state $x=\begin{bmatrix}x_{\mathrm{p}}&y_{\mathrm{p}}&\psi&v_{\mathrm{x}}&v_{\mathrm{y}}&\omega\end{bmatrix}^{\top}\in\mathbb{X}=\mathbb{R}^{6}$. The parameters of the considered system (6.1) are stated in Table 1 and represent a miniature race car from [70]. The Pacejka parameters $D_{\mathrm{f}}$ and $D_{\mathrm{r}}$ are considered to be unknown, i.e., $\theta=\begin{bmatrix}D_{\mathrm{f}}&D_{\mathrm{r}}\end{bmatrix}^{\top}\in\Theta=\mathbb{R}^{2}$, with given prior $\bar{\theta}_{0}=2\cdot\theta$. The system dynamics (6.1) are discretized using Euler forward with sampling time $\Delta t=10ms$, and the resulting discrete-time model is subject to uniformly distributed additive process noise $w_{\mathrm{x}}\in\mathbb{W}_{\mathrm{x}}=\{w_{\mathrm{x}}\in\mathbb{R}^{6}|\|w_{\mathrm{x}}\|_{\infty}\leq 0.01\}$.
At each time step, we obtain measurements of the position ($x_{\mathrm{p}},y_{\mathrm{p}}$) and yaw angle $\psi$ of the car, i.e.,
$y=\begin{bmatrix}x_{\mathrm{p}}&y_{\mathrm{p}}&\psi\end{bmatrix}^{\top}+w_{\mathrm{y}},$ (32)
which are subject to uniformly distributed additive measurement noise $w_{\mathrm{y}}\in\mathbb{W}_{\mathrm{y}}=\{w_{\mathrm{y}}\in\mathbb{R}^{3}|\ |w_{\mathrm{y}}|\leq\begin{bmatrix}0.2&0.2&0.01\end{bmatrix}^{\top}\}$. Such measurements can be provided by GPS in an outdoor environment or camera-based systems in an indoor environment, compare, e.g., [70]. The overall process and measurement noise is denoted as $w=\begin{bmatrix}w_{\mathrm{x}}^{\top}&w_{\mathrm{y}}^{\top}\end{bmatrix}^{\top}\in\mathbb{W}=\mathbb{W}_{\mathrm{x}}\times\mathbb{W}_{\mathrm{y}}\subseteq\mathbb{R}^{9}$.
Note that the augmented state $\begin{bmatrix}x^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is not uniformly detectable as the friction parameters cannot be estimated for low/zero slip $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$. In particular, if the car is driving straight ($\omega=0$, $\delta=0$, $v_{\mathrm{y}}=0$), the slip angles (31a)-(31b) and tire forces (31c)-(31d) are zero, i.e., no information about $\theta$ can be inferred. Consequently, no robustly stable estimator for the augmented state exists. On the other hand, the system (6.1)-(32) is in the form of (29) and hence the arguments in Section 5.5 apply, i.e., the derived stability results are applicable.
6.2 Numerical Estimation Results
Figure 4: Miniature race car example (Section 6): The top three plots show the states $x_{\mathrm{p}},\ y_{\mathrm{p}},\ \psi$ with corresponding sensor measurements and estimates resulting from a standard MHE approach with prior $\hat{\theta}_{t-M_{t}}$, the MHE approach without adaptation and fixed parameter value $\bar{\theta}_{0}$, and our proposed MHE approach with prior $\bar{\theta}_{0}$. The bottom three subplots show the estimation error for $v_{\mathrm{x}},\ v_{\mathrm{y}},\ \omega$ resulting from all three MHE approaches (dotted line) and a moving average thereof over $40$ time steps (solid line).
Figure 5: Miniature race car example (Section 6): The top three subplots show the normalized parameter estimation errors resulting from a standard MHE approach with prior $\hat{\theta}_{t-M_{t}}$, the MHE approach without adaptation and fixed parameter value $\bar{\theta}_{0}$, and our proposed MHE approach with prior $\bar{\theta}_{0}$ (dotted line) and a moving average thereof over $40$ time steps (solid line). The bottom two subplots show the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ and the corresponding estimates thereof resulting from all three MHE approaches.
In the following, we consider a scenario where the car is driving straight ($u_{t}=\begin{bmatrix}0&0.01\end{bmatrix}^{\top}$), takes a right turn between time $3s$ and $6s$ ($u_{t}=\begin{bmatrix}-0.03&0.01\end{bmatrix}^{\top}$), and drives straight again, see the top three subplots of Figure 4. We compare the following three approaches: (i) the proposed MHE approach with prior weighting w.r.t $\bar{\theta}_{0}$ (Section 4), (ii) a standard MHE approach for state and parameter estimation using updated prior $\hat{\theta}_{t-M_{t}}$ (Section 2.3), and (iii) an MHE approach for state estimation without parameter adaptation, i.e., the proposed MHE formulation but with fixed parameter estimate $\hat{\theta}_{t}=\bar{\theta}_{0}$. All three estimators are implemented in Python using CasADi [72] and the solver IPOPT [73]. The code will be made available online. We choose $Q=\mathrm{diag}(10^{4}\cdot I_{6},25\cdot I_{2},10^{4})$ and $R=\mathrm{diag}(25\cdot I_{2},10^{4})$, reflecting the bounds in $\mathbb{W}$. Furthermore, we choose $\eta=0.9$, $P_{\mathrm{x}}=P_{2,\mathrm{x}}=I_{6}$, $P_{\theta}=P_{2,\theta}=I_{2}$, a horizon $M=20$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. The system is initialized with $x_{0}=\begin{bmatrix}0&0&0&4&0&0\end{bmatrix}^{\top}$, while the initial estimate is chosen as $\bar{x}_{0}=\begin{bmatrix}0&0&0&3.8&0&0\end{bmatrix}^{\top}$.
Figure 4 shows the ground truth states and state estimation errors resulting from all three MHE approaches. In Figure 5, the corresponding estimation errors of the unknown parameters $D_{\mathrm{r}}$ and $D_{\mathrm{f}}$, and the estimates of the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ are shown. The first three states ($x_{\mathrm{p}}$, $y_{\mathrm{p}}$, and $\psi$) are accurately estimated from the noisy measurements. The proposed MHE approach results in smaller estimation errors for the longitudinal, lateral, and angular velocities, $v_{\mathrm{x}}$, $v_{\mathrm{y}}$, $\omega$, respectively, sometimes over one order of magnitude compared to the two other MHE schemes. The MHE approach with prior $\hat{\theta}_{t-M_{t}}$ leads to drifting parameter estimates while the car is driving straight. When the car is turning, this yields bad estimates of the tire forces and consequently large state estimation errors. The MHE approach without adaptation leads to inaccurate estimates of the tire forces during the turn, which also results in increased state estimation erros. In contrast, the proposed MHE approach prevents parameter drift when driving straight, while the parameter estimates are improved during the turn and consequently lead to accurate estimates of the tire forces and reduced state estimation errors.
The considered problem of jointly estimating state and friction parameters in a car is of high practical relevance, since friction parameters are difficult to identify and often varying, e.g., due to aging of tires or changes in the road conditions. The problem is frequently addressed using MHE, see, e.g., [5, 6]. The presented numerical results show that a standard MHE approach for joint state and parameter estimation can lead to large state and parameter estimation errors. This can have detrimental effects, especially if the resulting state estimate is used to compute control inputs to a system. As shown, these issues can be overcome by a very simple design change: The regularization w.r.t. a constant prior $\bar{\theta}_{0}$ instead of a parameter estimate $\hat{\theta}_{t-M_{t}}$ obtained online. While the proposed regularization w.r.t. a constant prior $\bar{\theta}_{0}$ ensures robustness of the resulting state estimates, it still allows for sufficient adaptation and consequently improved state estimates compared to an MHE with no parameter adaptation at all.
7 Conclusion
State estimation algorithms require some form of online adaptation to address nonlinear systems with unknown or time-varying parameters. In this paper, we showed that standard moving horizon estimation (MHE) for joint state and parameter estimation can fail in the absence of persistency of excitation (PE), e.g., if a system is well regulated or over-parametrized. We proposed a novel MHE formulation with regularization on a (constant) prior estimate of the unknown system parameter and established practical stability of the resulting state estimate independent of PE. We discussed how many existing results from MHE, adaptive estimation, and functional estimation relate to the proposed MHE approach. Finally, we showed practical applicability using a car example.
While the proposed regularization on a given prior parameter estimate allows us to establish stability irrespective of PE, it may deteriorate the estimation performance compared to a standard MHE approach in case PE is sufficiently large. An interesting question for future research is to extend the proposed MHE approach to achieve an estimation performance comparable to a standard MHE scheme in case of large PE, while the robust stability properties irrespective of PE of the proposed approach are preserved.